OpenStudy (safari321):

what is -square root 100 + 8 square root 10

OpenStudy (turingtest):

\[-\sqrt{100}+8\sqrt{10}\]is the question?

OpenStudy (safari321):

yes

jhonyy9 (jhonyy9):

is -10+8sqrt10

OpenStudy (turingtest):

\[-\sqrt{100}=-10\] so\[-10+8\sqrt{10}\]

OpenStudy (safari321):

is that the answer

OpenStudy (turingtest):

that's simplified, if that's what you mean by "answer".

OpenStudy (safari321):

okay

OpenStudy (turingtest):

15.29822128.... if you wanted a number but that's what god made calculators for :)

OpenStudy (safari321):

okay thank you